OpenStudy (anonymous):

Find the three arithmetic means between 5 and 17.

OpenStudy (anonymous):

d = difference between the means . 5 + d, 5 + 2d, and 5 + 3d. Since the numbers are equally spaced from 17, we require 5 + 4d = 17. d = 3. :) 5 +3, 5+2(3), and 5 + 3(3)

OpenStudy (anonymous):

If we let d be the difference between the arithmetic means, then we have that the arithmetic means are: 5 + d, 5 + 2d, and 5 + 3d. However, since the numbers are equally spaced from 17, we require 5 + 4d = 17 <==> d = 3. Thus, the numbers are: 5 + 3 = 8, 5 + 6 = 11, and 5 + 9 = 14. I hope this helps!